### v1

@@ Abstract
Dynamic resource management is an increasingly important capability of High Performance Computing systems, as it enables jobs to adjust their resource allocation at runtime. This capability has been shown to reduce workload makespan, substantially decrease job waiting times and improve overall system utilization. In this context, malleability refers to the ability of applications to adapt to new resource allocations during execution. Although beneficial, malleability incurs significant reconfiguration costs, making the reduction of these costs an important research topic.
Some existing methods for MPI applications respawn the entire application, which is an expensive solution that avoids the reuse of original processes. Other MPI methods reuse them, but fail to fully release unneeded processes when shrinking, since some ranks within the same communicator remain active across nodes, preventing the application from returning those nodes to the system. This work overcomes both limitations by proposing a novel parallel spawning strategy, in which all processes cooperate in spawning before redistribution, thereby reducing execution time. Additionally, it removes shrinkage limitations, allowing better adaptation of parallel systems to workload and reducing their makespan. As a result, it preserves competitive expansion times with at most a $1.25\times$ overhead, while enabling fast shrink operations that reduce their cost by at least $20\times$. This strategy has been validated on both homogeneous and heterogeneous systems and can also be applied in shared-resource environments.
@@ 1 Introduction
To address this issue, Resource Management Systems (RMS) must be able to distribute resources dynamically, a capability known as Dynamic Resource Management (DRM). This capability enables RMSs to adjust job allocations at runtime, provided the applications are malleable, that is, capable of resizing and adapting to changes in resource availability during execution. Such flexibility allows applications to achieve more effective optimization of resource utilization [4], throughput [9], energy efficiency [3, 11], and I/O performance [23].
@@
Recent studies by the US Exascale Computing Project [2] suggest that DRM could significantly enhance the performance of both applications and systems, making it a key contribution for improving relevant metrics in future platforms [13]. However, its adoption remains limited compared to other MPI features [6], primarily due to the complexity involved in adapting applications to support malleability and enabling systems to dynamically reallocate resources at runtime.
@@
There are already libraries available to perform reconfigurations in applications; some of the most relevant are DMR[9], FlexMPI[14], MaM[16] or DPP[8]. A particularly challenging scenario for most of these tools is shrinking, i.e., reducing the number of active processes in a job at runtime to free resources or adapt to changing workload demands. This operation is limited by the rigid structure of MPI communicators, especially the standard MPI_COMM_WORLD (MCW), which does not allow the partial termination of ranks.
@@
As shown in [20], terminate-based (TB) methods can shrink without spawning new processes, providing a time efficient alternative to traditional spawn-based (SB) approaches. Nevertheless, a major limitation of TB shrinking is the persistence of zombie processes, ranks that remain asleep until the application terminates, as they cannot be fully removed from the global MCW. This creates an issue: the nodes on which these zombies are mapped cannot be returned to the RMS until they terminate, preventing node reuse. A common workaround involves isolating each MCW communicator on a separate node, allowing finer control over process groups. However, this solution depends on node-specific spawning, which suffers from poor scalability due to its inherently sequential nature [22].
To overcome these constraints, we propose a novel technique that retains the benefits of TB shrinking while addressing its key limitations, supporting both expansion and shrinkage at minimal cost. Our approach isolates each MCW communicator on a separate node and enables all available processes to perform spawning in parallel rather than sequentially, thereby reducing the cost of expansion. In addition, it avoids spawning new processes during shrinkage and eliminates the persistence of zombie processes while maintaining the high performance observed in TB methods. These improvements make our solution particularly suitable for DRM in large-scale HPC environments, with full support for heterogeneous systems and shared resources. The approach has been included in the MaM library [16], which integrates multiple techniques for resizing and data redistribution, allowing the selection of the optimal solution depending on the context.
The main contributions of this work are:
@@
  • The design and implementation of two novel parallel spawning algorithms tailored to the MPI model, enabling scalable malleability for HPC applications.
  • The extension of the MaM library [16] to support dynamic reconfiguration in heterogeneous queue systems and shared-resource environments.
  • A comprehensive performance evaluation of the proposed algorithms on both homogeneous and heterogeneous clusters, demonstrating their advantages over the previous best method in terms of scalability and reconfiguration overhead.
@@
2 Background
According to the classification proposed by Feitelson and Rudolph [5], applications executed in large-scale computing facilities can be grouped into four categories. This classification is based on two main criteria: (1) whether the number of processes can change during execution, i.e., whether reconfiguration is supported, and (2) who determines the job size, either the user or the RMS. Table 1 summarizes this classification.
@@ 2 Background
Rigid Job Static User Can only be executed with a fixed number of processes defined by the user. No reconfiguration is allowed.
@@
Evolving Job Dynamic User Includes a user-defined reconfiguration scheme. The RMS must meet the resource requests for the job to continue executing.
@@
    Process management. The RMS assigns a new number of cores to the job, which dictates how many MPI ranks should be created or terminated. Processes before resizing are considered sources, while those that continue after resizing are considered targets.
@@
Incorporating all these stages into a parallel application is a complex task that goes far beyond simple integration with the RMS. It requires modifications to the application’s structure to support dynamic creation and termination of processes, manage data redistribution efficiently, and reconfigure communication patterns on the fly. These challenges make dynamic applications difficult to implement from scratch, often requiring dedicated tools and frameworks that simplify the transformation of traditional applications into dynamic ones. Without such support, adapting applications to handle resizing becomes both error-prone and time-consuming.
@@
One malleable solution is Flex-MPI [14], a library built on top of MPICH [21] that bases its spawn operations on the MPI_Comm_spawn function. Its primary goal is to help applications become dynamic and the main result of this solution is the Merge method, which is described below. When an application expands from $x$ processes to $y$ processes, only the additional yxy-x processes are spawned and added to the existing ones. Then, each new process is created through a separate call to the spawn function, which allows a fine grain when shrinking the application, as dynamic processes can be terminated at any time. Nevertheless, there is a limitation with the minimum number of ranks: the initial number of ranks cannot be terminated until the end of the application.
@@
A similar solution is MaM [16], a malleability module also built on top of MPICH and within the MPI scope, developed to analyze and compare different malleability techniques. Like Flex-MPI, it bases its spawn operation on the MPI_Comm_spawn function, providing a Merge method, but without the limitation of the minimum number of processes. MaM is part of the Proteo framework [19], which helps transform applications into malleable ones, supports the generation of dynamic workloads, and enables performance analysis of different resize operations to select the optimal alternative for a given situation.
@@
Another approach is DPP [8], which explores the new concept of MPI Sessions, extending it to implement on-the-fly dynamic mechanisms. Since the sessions model operates outside the traditional MPI environment, the MCW communicator does not exist and some functions, such as MPI_Comm_spawn, cannot be used. However, alternative non-standard MPI functions are provided to allow communication between ranks. To spawn processes, it uses a custom dynamic RMS, which determines whether a job should get or return resources. When a job gets additional resources, it informs a PMIx Reference Runtime Environment (PRRTE) daemon, which then spawns of new processes. Once the processes are ready, the job is informed of the new resources and how to connect with them, which is managed by DPP. During shrink, processes do not have the limitations of the traditional environment and can be terminated normally.
@@
Lastly, an alternative solution is Dynamic Management of Resources library (DMRlib) [9], which consists of two components: a parallel distributed runtime based on MPI and an extension of SLURM to enable the execution of malleable jobs. Initially, the library only allowed the simplest approach, the Baseline method, which always spawns the new number of processes $y$ and terminates the previous $x$ processes. However, the library now supports other environments for process management, such as DPP [7] or MaM [10], providing a wider range of possibilities.
@@
For further details on MPI spawn based malleability, see [1], while for a broader review of dynamic resource management [24] is recommended.
@@ 3 MaM overview
MaM is the Proteo module responsible for completing malleability. It provides a set of features designed to transform traditional MPI applications into malleable ones, requiring minimal effort from developers. Additionally, it defines a simple and intuitive interface, which lowers the barrier to integrating malleability into existing parallel applications [16].
@@
Specifically, MaM implements multiple methods and strategies to handle the most critical stages of malleability, process management and data redistribution, allowing applications to dynamically adjust their number of MPI ranks during execution.
@@
For the process management stage, i.e., resizing from an initial group of NS processes to a new group of NT processes, MaM offers two main approaches: the Baseline method, which always creates a new group of processes, and the Merge method, which reuses existing processes and only spawns or terminates the difference in ranks. These methods can be combined with two strategies: the Asynchronous strategy, which overlaps process creation with application execution to reduce downtime, and the Single strategy, where only one process performs the spawn operation and informs the rest afterwards. This flexible design allows users to decide the preferred configuration according to their application needs.
For the data redistribution stage, MaM distinguishes between constant and variable data, as this affects how and when the redistribution can take place. While variable data requires execution to be paused until the transfer is complete, constant data can be redistributed asynchronously, allowing overlap with ongoing computation. The library also supports three main communication methods for this stage: Point-to-Point, Collective, or One Sided primitives.
Additionally, asynchronous redistribution can be implemented through two strategies: Non-blocking communication, which relies on routines like MPI_Test to monitor progress, and Threading, where dedicated threads handle data transfer independently of the main computation.
Using MaM to incorporate malleability into an application requires specifying a method with none, one, or more of the strategies for stages 22 and $3$. A more detailed description of all alternatives related to process management can be found in [20] and to automatic data redistribution in [15] and [18].
4 Methods for spawning processes in parallel
@@
As previously discussed, the MaM shrinkage approach is constrained by its inability to fully terminate unneeded processes during execution. As a consequence, is not possible to return those resources until the MCW they pertain terminates completely. To overcome these issues, this work proposes new strategies based on parallel spawning, which allows processes to be resized efficiently and overcome said situation.
@@
Parallel process creation is inherently complex and is coordinated across multiple phases: (1) It begins with the spawn of new process groups and opening communication ports. (2) it is followed by a synchronization phase, where all groups wait until every required port is ready. (3) Next, the newly spawned groups are merged into a single communicator through the ports. (4) Finally, a global reordering of ranks is performed to maintain a logical order between nodes.
@@
The first phase has been examined through two alternatives: A simpler option only for homogenous systems called Hypercube strategy, and an improved but more complex one that works for heterogenous or shared systems, called Iterative Diffusive strategy. The following sections detail each of these steps, highlighting the specific operations involved.
4.1 Hypercube: Parallel Spawning for Homogeneous Systems
@@
This parallel strategy is based on performing a call to the MPI_Comm_spawn function for each node that is to be used, assigning all available cores of the node to each newly created process.
@@
The procedure is organized into a series of steps. In each step, all active processes, whether source or target, concurrently execute a spawn operation, each one acting on a different node. As the number of target processes increases, so does the number of processes participating in the spawning operation. Consequently, with each successive step, a growing number of processes are involved in initiating new spawns, enabling a scalable and distributed expansion across nodes.
@@
In the first step, only the source processes participate in the operation. In total, up to $C\cdot I$ processes can be launched, where $C$ is the number of cores per node and $I$ is the number of initial nodes. From the second step onward, both the original and previously created processes participate, exponentially increasing the spawning capacity. However, in the final step, only the first $x$ processes participate, where $x$ corresponds to the remaining number of nodes needed to reach the desired number of nodes.
@@
As an example, consider a system with 20 cores per node. Starting with a single fully occupied node, the first step could launch up to 20 additional nodes. In the second step, with 420 processes available (21 nodes with 20 cores each), it would be possible to occupy 410 additional nodes by creating 20 processes per node.
@@
From a technical standpoint, this strategy is implemented using MPI_Comm_spawn in combination with the MPI_COMM_SELF communicator and an MPI_Info object that specifies the target node for each new group of processes. Furthermore, all processes in each node are assigned the group identifier ranging from 0 to the total number of groups to be created (group_id), identifying which group it belongs.
The number of steps $s$ required to reach a total of $N$ nodes, starting from $I$ initial nodes and with $C$ cores per node, can be calculated using Equation 1. $N$ and $I$ are calculated as $\frac{NT}{C}$ and $\frac{NS}{C}$, respectively, where $NT$ is the amount of targets, and $NS$ the amount of sources.
${s}=\frac{\ln(N/I)}{\ln(1+C)}$ (1)
From this, the total number of nodes $T_{s}$ at the end of step $s$ can be computed as:
${T_{s}}=(1+C)^{s}\cdot I$ (2)
And the corresponding number of processes spawned $t_{s}$ in step $s$ is simply:
${t_{s}}=T_{s}\cdot C$ (3)
@@
Figure 1 illustrates this strategy with a simple example in which the number of processes is expanded from $NS=1$ to $NT=8$, assuming an architecture with $C=1$ core per node. In this case, seven additional groups are created over three steps. The process is visualized as a cube, where the edges indicate which group spawns which in each step, and the vertices are labelled with the corresponding group_id, or with the label $I$ for the initial group. By the end of the procedure, a total of eight nodes are occupied.
@@
Figure 1: Generation of 7 groups using parallel spawning. Each step is represented with different markers.
4.2 Iterative Diffusive: Recursive Algorithm for Heterogeneous Systems
This method improves upon the previous scheme by including the ability to create process groups of variable sizes at each step, which the Hypercube is not capable of. This makes it particularly well-suited to heterogeneous or shared computing environments, where nodes may have different capacities or resources may be distributed unevenly.

### v2

@@ Abstract
Dynamic resource management is an increasingly important capability of High Performance Computing systems, as it enables jobs to adjust their resource allocation at runtime. This capability can reduce workload makespan, substantially decreasing job waiting times and optimizing resource allocation. In this context, malleability refers to the ability of applications to adapt to new resource allocations during execution. Although beneficial, malleability incurs significant reconfiguration costs, making the reduction of these costs an important research topic. Some existing solutions for MPI applications respawn the entire application, which is an expensive solution that avoids the reuse of original processes. Other MPI solutions reuse them, but fail to fully release unneeded processes when shrinking, since some ranks within the same communicator remain active across nodes, preventing the application from returning those nodes to the system. This work overcomes both limitations by proposing a novel parallel spawning strategy, in which all processes cooperate in the spawning. This allows expansions to reuse processes while also terminating unneeded ones. This strategy has been validated on two systems with either machines with equal or different numbers of cores. Experiments show that this strategy preserves competitive expansion times with at most a $1.13\times$ and $1.25\times$ overhead for equal and different number of cores per node, respectively. More importantly, it enables fast shrink operations that reduce their cost by at least $1387\times$ and $20\times$ in the same scenarios.
@@
Dynamic Resource Management , Malleability , MPI , Process Management , Heterogenous Allocations
journal: Journal of Parallel and Distributed Computing
{graphicalabstract}[Uncaptioned image]
@@
[1]organization=Dpto. de Ingeniería y Ciencia de los Computadores, Universitat Jaume I, addressline=Av. Vicent Sos Baynat, s/n , city=Castelló, postcode=12071, state=Comunitat Valenciana, country=Spain
@@ 1 Introduction
To address this issue, Resource Manager System (RMS) must be able to distribute resources dynamically, a capability known as Dynamic Resource Management (DRM). This capability enables RMS s to adjust job allocations at runtime, provided that the applications are malleable, that is, capable of resizing and adapting to changes in resource availability during execution. Such flexibility allows applications to achieve more effective optimization of resource utilization [4], throughput [11], energy efficiency [3, 10], and I/O performance [21].
@@
A prior study by the US Exascale Computing Project [2] suggest that DRM could significantly enhance the performance of both applications and systems, making it a key contribution for improving relevant metrics in future platforms [13]. However, its adoption remains limited compared to other MPI features [6], primarily due to the complexity involved in adapting applications to support malleability and enabling systems to dynamically reallocate resources at runtime.
@@
There are already libraries available to perform reconfigurations in applications; some of the most relevant are DPP [8], DMR [11], FlexMPI[16], or MaM [14]. A particularly challenging scenario for most of these tools is shrinking, i.e., reducing the number of active processes in a job at runtime to free resources or adapt to changing workload demands. This operation is limited by the rigid structure of the standard MPI communicator MPI_COMM_WORLD (MCW), which does not support the termination of a subset of ranks, even if separated with a MPI_Comm_split function.
@@
In the paper [17] two methods, Baseline and Merge, are evaluated with different strategies. The latter method allowed to shrink the application by the conversion of processes into zombies, which we denote as Zombies Shrinkage (ZS). The ZS offered lower reconfiguration times than traditional Spawn Shrinkage (SS) obtained with the Baseline method. Nevertheless, a major limitation of ZS is the persistence of zombie processes, ranks that remain asleep until the application terminates, as they cannot be fully removed from the global MCW. This creates an issue: the nodes on which these zombies are mapped cannot be returned to the RMS until they terminate, preventing node reuse. A common workaround involves isolating each MCW communicator on a separate node, allowing finer control over process groups. However, this solution depends on node-specific spawning, which suffers from poor scalability due to its inherently sequential nature [20]. Yet, by using this technique it is possible to shrink the application by terminating those processes, an approach that we denote as Termination Shrinkage (TS), which overcomes the key limitation of ZS, its inability to fully terminate processes and make them persist as zombies.
To overcome the scalability constraints identified in [20], we propose a novel technique that preserves scalability during process spawning while enabling the use of TS for shrinkages. Our technique also isolates each MCW communicator on a separate node; however, instead of performing this step sequentially, it is carried out in parallel, allowing each available process to initiate an independent spawn. Additionally, newly spawned processes participate in subsequent spawn operations within the same reconfiguration if further spawning is required.
These improvements make our solution particularly suitable for DRM in large-scale HPC environments, with full support for nodes with different numbers of cores, and for jobs with different numbers of cores on each node. Consequently, to evaluate the novel technique, this work considers two types of node allocations: homogeneous, where jobs use the same number of cores in each node, and heterogeneous, where the number of cores used in each node can differ. Additionally, the improvements can be used for oversubscription in both types of allocation, where the number of processes exceeds the number of cores allocated to the job. Nevertheless, it is not considered the challenge of load balancing between different types of hardware, which is postponed as future work.
This approach has been included in the MaM library [14], which integrates multiple techniques for resizing and data redistribution, allowing the selection of the optimal solution depending on the context. The main contributions of this work are:
@@
  • The design and implementation of a novel parallel spawning algorithm tailored to the MPI model, enabling scalable malleable executions for HPC applications in homogeneous node allocations.
  • An adaptation of the previous technique to enable resizing in heterogeneous node allocations.
  • The capability to shrink malleable applications by terminating processes (TS) without spawning new ones (SS) or leaving them as zombies (ZS).
  • A comprehensive performance evaluation of the proposed algorithms on jobs with homogeneous node allocations, demonstrating their advantages.
  • A performance evaluation of the extended algorithm for jobs with heterogeneous node allocations, demonstrating its advantages.
@@
2 Related Work
According to the classification proposed by Feitelson and Rudolph [5], applications executed in large-scale computing facilities can be grouped into four categories (see Table 1). This classification is based on two main criteria: (1) whether the number of processes can change during execution, i.e., whether reconfiguration is supported, and (2) who determines the new job size, either the application or the RMS.
@@ 2 Related Work
Rigid Job Static Application Can only be executed with a fixed number of processes defined by the user. No reconfiguration is allowed.
@@
Evolving Job Dynamic Application Includes a user-defined reconfiguration scheme. The job will only continue if the RMS satisfies the resource request; otherwise, it waits or it is cancelled.
@@
    Process management. The RMS assigns a new number of cores to the job, which determines how many MPI ranks should be created or terminated. Processes before resizing are considered sources, while those that continue after resizing are considered targets.
@@
Incorporating all these stages into a parallel application is a complex task that goes far beyond simple integration with the RMS. It requires modifications to the application’s structure to support dynamic creation and termination of processes, the efficient management of data redistribution, and the reconfiguration of communication patterns on the fly. These challenges make dynamic applications difficult to be implemented from scratch, since they often require dedicated tools and frameworks to simplify the transformation of traditional applications into dynamic ones. Without such support, adapting applications to handle resizing becomes both error-prone and time-consuming.
@@
One malleable solution is Flex-MPI [16], a library built on top of MPICH [19] that bases its spawn operations on the MPI_Comm_spawn function. The primary goal of this solution is to help applications become dynamic and the main method to resize in Flex-MPI is the Merge method, which is described below. When an application expands from $x$ processes to $y$ processes, only the additional yxy-x processes are spawned and added to the existing ones. Then, each new process is created through a separate call to the spawn function, which allows a fine grain when shrinking the application, as dynamic processes can be terminated at any time. Nevertheless, there is a limitation with the minimum number of ranks: the initial number of ranks cannot be terminated until the end of the application.
@@
A similar solution is MaM [14], a malleability module also built on top of MPICH and within the MPI scope, developed to analyze and compare different malleability techniques. Like Flex-MPI, it bases its spawn operation on the MPI_Comm_spawn function and implements Baseline and Merge methods, in which the former always spawns the new number of processes $y$ and terminates the previous $x$ processes. MaM is part of the Proteo framework [18], which helps transform applications into malleable ones. It supports the generation of dynamic workloads, and enables performance analysis of different resize operations to select the optimal alternative for a given situation, by simulating different scenarios.
@@
Another approach is DPP [8], which explores the new concept of MPI Sessions, extending it to implement on-the-fly dynamic mechanisms. Since the sessions model operates outside the traditional MPI environment, the MCW communicator does not exist and some functions, such as MPI_Comm_spawn, cannot be used. However, alternative non-standard MPI functions are provided to allow communication between ranks. DPP uses a custom dynamic RMS, which determines whether a job should get or return resources. When a job gets additional resources, it notifies a PMIx Reference RunTime Environment (PRRTE) daemon, which then spawns the new processes. Once the processes are ready, the job is notified of the new resources and how to connect with them, which is managed by DPP. During shrink, processes do not have the limitations of the traditional environment and can be terminated arbitrarily.
@@
Lastly, an alternative solution is Dynamic Management of Resources library (DMRlib) [11], which consists of two components: a parallel distributed runtime based on MPI and an extension of Slurm to enable the execution of malleable jobs. Initially, the library only allowed the simplest approach, the Baseline method. However, the library now supports other environments for process management, such as MaM [9] or DPP [7], providing a wider range of possibilities.
@@
For further details on MPI spawn based malleability, see [1], while a broader review of dynamic resource management can be found in [22].
@@ 3 MaM overview
MaM is the Proteo module responsible for completing malleability. It provides a set of features designed to transform traditional MPI applications into malleable ones, requiring minimal effort from developers. Additionally, it defines a simple and intuitive interface, which makes easier to integrate malleability into existing parallel applications [14].
@@
Specifically, MaM implements multiple methods and strategies to handle the most critical stages of malleability, process management and data redistribution. These techniques allow applications to dynamically adjust the number of MPI ranks during execution. Using MaM to incorporate malleability into an application requires specifying a method with none, one, or more of the strategies for process management and data redistribution.
@@
For the process management stage, i.e., resizing from an initial group of Source processes (NS) to a new group of Target processes (NT), MaM offers two main methods: the Baseline method, which always creates a new group of processes, and the Merge method, which reuses existing processes and only spawns or terminates the difference in ranks. These methods can be combined with two strategies: the Asynchronous strategy, which overlaps process creation with application execution to reduce downtime, and the Single strategy, where only one process performs the spawn operation and informs the rest afterwards. This flexible design allows users to decide the preferred configuration according to their application needs.
4 Strategies for spawning processes in parallel
@@
As previously discussed, the ZS approach is constrained by its inability to fully terminate unnecessary processes during execution. As a consequence, it is not possible to return the corresponding nodes until MCW, to which they belong, has terminated completely. However, a previous study [17] using a test system showed for shrinkages that the ZS method delivered a speedup of $36\times$ over SS methods. Consequently, reaching a solution for shrinkage with TS methods would prove a valuable advantage to reduce the overhead caused by resizing. To overcome these challenges, this work proposes new strategies based on parallel spawning, which allows processes to be resized efficiently and create an MCW for each node, facilitating the use of a TS approach.
@@
Parallel process creation is inherently complex and is coordinated across multiple phases: (1) New processes groups are spawned and communication ports are opened. (2) This is followed by a synchronization phase, where all groups wait until every required port is ready. (3) Next, the newly spawned groups are merged into a single communicator through the ports. (4) Finally, a global reordering of ranks is performed to maintain a logical order between nodes.
@@
The first phase has been addressed through two alternative approaches: A simpler option only for homogeneous node allocations called Hypercube strategy, and an improved but more complex one that also works for jobs with heterogeneous node allocations, called Iterative Diffusive strategy. The following subsections detail these approaches, and explain how the four phases are optimized, highlighting the specific operations involved in each step.
4.1 Hypercube Strategy
@@
This parallel strategy is based on invoking the MPI_Comm_spawn function once for each node to be used (up to the total number of required MPI ranks), assigning all available cores of the node (process group) to each newly spawned process.
@@
The procedure is organized into a series of steps. In each step, all existing processes concurrently execute a spawn operation, each one acting on a different node. As the number of spawned processes increases, so does the number of processes participating in the spawning operation. Consequently, with each successive step, a growing number of processes are involved in initiating new spawns, enabling a scalable and distributed expansion across nodes.
@@
In the first step, only the source processes participate in the operation. In total, up to $C\cdot I$ processes can be launched, where $C$ is the number of cores per node and $I$ is the number of initial nodes. From the second step onward, both the original and previously spawned processes participate, exponentially increasing the spawning capacity. However, in the final step, only the first $x$ processes participate, where $x$ corresponds to the remaining number of nodes needed to reach the desired number of nodes.
@@
As an example, consider a system with $20$ cores per node. Starting with a single fully occupied node, the first step could launch up to $20$ additional nodes. In the second step, with $420$ processes available ($21$ nodes with $20$ cores each), it would be possible to occupy $420$ additional nodes, creating $20$ processes per node.
@@
From a technical standpoint, this strategy uses MPI_Comm_spawn in combination with the MPI_COMM_SELF communicator and a MPI_Info object that specifies the target node for each new group of processes. Furthermore, all processes in each node are assigned a group identifier ranging from 0 to the total number of groups to be spawned (group_id), identifying which group it belongs.
From a mathematical perspective, the strategy can be described as a geometric growth, where each existing process expands into $C$ new nodes while the process that creates each node remains active. Therefore, the effective growth factor per step is $C+1$. The initial number of nodes, $I$, is determined by the total number of source processes NS as:
$I=\frac{NS}{C}\quad,\quad NS\;mod\;C=0.$
Using this initial condition and the growth factor, the total number of nodes at step $s$ is given by:
$T_{s}=\begin{cases}(C+1)^{s}\cdot I-I,&\text{if Method}=\text{Baseline}\\ (C+1)^{s}\cdot I,&\text{if Method}=\text{Merge}\end{cases}\quad,$ (1)
where $T_{s}$ represents the total nodes at step $s$ and $C+1$ is the growth factor including the cores and the process that spawns them. Additionally, it considers the chosen spawning method, either Baseline or Merge. For the former, the initial nodes ($I$) are not considered to reach the target nodes ($N$), as the source processes will be removed after the reconfiguration. Thus, the corresponding total number of processes at step $s$ is simply obtained by multiplying the number of nodes by the cores per node:
${t_{s}}=C\cdot T_{s}.$ (2)
Finally, the total number of steps required to reach a target of $N$ nodes can be derived from Equation 1 by solving for $s$ as:
$s=\left\lceil\frac{\ln(N/I)}{\ln(C+1)}\right\rceil,$ (3)
where, $N$ represents the desired total number of nodes, calculated as:
$N=\frac{NT}{C}\quad,\quad NT\;mod\;C=0,$
where, NT is the total number of target processes.
@@
Figure 1 illustrates this strategy with a simple example in which the number of processes is expanded from $NS=1$ to $NT=8$, assuming an architecture with $C=1$ core per node. In this case, seven additional groups are spawned over the course of three steps. The process is visualized as a cube, where the edges indicate which group spawns which in each step, and the vertices are labeled with the corresponding group_id, or with the label $I$ for the initial group. By the end of the procedure, a total of eight nodes are occupied.
@@
Figure 1: Generation of $7$ groups using parallel spawning. Each step is represented with different markers, and the number marks the group identifier.
This strategy removes the shared MCW when expanding across multiple nodes. Therefore, it enables shrinkage via the TS method by releasing the excess processes.
4.2 Iterative Diffusive Strategy
The main limitation of the Hypercube strategy is that all spawned process groups have a constant size. The Iterative Diffusive parallel strategy can use variable sizes for each process group. This makes it particularly well-suited for jobs with heterogeneous node allocations.